OpenStudy (anonymous):

Can someone help me with Algebra 2 functions?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Let f(x) = x2 - 8x + 5. Find f(-1). -4 -3 13 14 Question 2 (Multiple Choice Worth 1 points) Let f(x) = 2x - 6 Solve f-1(x) when x = 2. -5 4 -2 7 Question 3 (Multiple Choice Worth 1 points) Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 1 3 4 10 Question 4 (Multiple Choice Worth 1 points) Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). -7 -3 10 33 Question 5 (Multiple Choice Worth 1 points) Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). -21 -3 3 21 Question 6 (Multiple Choice Worth 1 points) Let f(x) = thirty over the quantity of 3 x minus Find f(-2). -6 -3 3 6 Question 7 (Multiple Choice Worth 1 points) Let f(x) = 5x + 12. Find f-1(x). 12x + 5 1 over quantity 5 x plus 12 5 over quantity x minus 12 quantity of x minus 12 over 5 Question 8 (Multiple Choice Worth 1 points) Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27 Question 9 (Multiple Choice Worth 1 points) Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (anonymous):

Question 10 (Multiple Choice Worth 1 points) Let f(x) = x2 - 81. Find f-1(x). �square root of the quantity x plus 81 9square root of x 1 over quantity x squared minus 81 x squared over 81

OpenStudy (mertsj):

Replace x with -1 in the first one.